Find the cylindrical coordinates of the points where the cartesian coordinates are i)(3,3,4) ii) (√5,1,2)


1
Expert's answer
2022-05-03T07:07:53-0400

"r^2=x^2+y^2"

"tan \phi=y/x"

z=z

1. x=3, y=3, z=4

"r= \sqrt{3^2+3^2}= \sqrt{18}"

"tan \phi =1"

"\phi =\pi/4"

z=4

So, the cylindric coordinates are ("\sqrt{18},\pi/4,4)"

2. "x=\sqrt{5}" ,y=1, z=2

"r=\sqrt{\sqrt{5}^2+1^2}=\sqrt6"

"tan\phi=1/\sqrt{5}"

"\phi=\pi/7.5"

z=2

So, the cylindric coordinates are ("\sqrt6, \pi/7.5,2)"